OpenStudy (anonymous):

-15c + 7c + 1 = 3 - 8c

OpenStudy (lgbasallote):

first..combine -15c and +7c what do you get?

OpenStudy (anonymous):

-8c

OpenStudy (lgbasallote):

right. so you have -8c + 1 = 3 - 8c now add 8c to both sides..what do you get?

OpenStudy (anonymous):

8c + 1 = 3 - 8c ?

OpenStudy (lgbasallote):

nope. you have -8c + 1 = 3 - 8c if you add 8c to both sides you get -8c + 8c + 1 = 3 - 8c + 8c right?

OpenStudy (anonymous):

ohh... sorta get it.

OpenStudy (lgbasallote):

wonderful. so now..combine the like terms

OpenStudy (theviper):

-15c + 7c + 1 = 3 - 8c -8c+1=3-8c \(1\neq3\) No answer.

OpenStudy (lgbasallote):

i have to go now so i hope you can go on from where i stopped @Rubie